A rectangular carpet has a perimeter of 232 inches. The length of the carpet is 26 inches more than the width, what are the dimensions of the carpet
Answer by meli10494(1)

use the formula p=2L+2W
let p=232 in
let L=26+W
232=2(26+W)+2W plug the numbers in
232=52+2W+2W multiply and/or add
-52=-52+4W substract 52 from both sides
180=4W divide both sides by 4
45=w or w=45
then plug-in your answer in (26+W)
26+45= 71
so W=45 L=71